AN ASSESSMENT OF TERAHERTZ-FREQUENCY REMOTE-SENSING FOR APPLICATION TO A BIO-LOGICAL WARFARE ATTACK SCENARIO

\underline{D. L. WOOLARD}, Research Laboratory, Army Research Office, Research Triangle Park, NC; E. R. BROWN, University of California at Los Angeles, Los Angeles, CA; A. C. SAMUELS, J. O. JENSEN, Edgewood Chemical and Biological Center, Aberdeen Proving Ground, MD; T. GLOBUS, B. GELMONT, University of Virginia, Charlottesville, VA; M. WOLSKI, Naval Surface Warfare Center, Dahlgren, VA.

This presentation gives a detailed assessment of terahertz-frequency spectroscopy as a technique for the remote detection of biological warfare agents. Design studies are presented for a differential-absorption-radar (DAR) approach that utilizes the spectral signatures of Bacillus (B.) subtillus spores within the terahertz (THz) regime as the detection mechanism. The signature data used in these studies is taken from laboratory measurements performed on uniform thin films of B. Subtillus spores and the system performance is assessed for both incoherent and coherent detector modalities. These studies consider DAR remote sensing of biological (bio) clouds at significant ranges (i.e., 1 km) and include the effects of realistic atmospheric conditions. A high-level remote-sensor design is used to estimate the probabilities of detection (p_{d}) and false-alarm (p_{fa}) associated with this general technique. These studies suggest useful remote-detection performance can be achieved (i.e., p_{d} > 0.9 & p_{fa} ≪ 10^{−4} for bio-cloud densities < 103 cm^{−3}) at 1 km ranges if the THz signature information remains predictably stable under varying atmospheric conditions (e.g., changes in humidity, spore activity state, etc). Furthermore, a realistic bio-agent airframe attack scenario is utilized to demonstrate standoff detection of bio-clouds with ≈100% confidence while outside the threat-level concentrations. All together, these results demonstrate that standoff detection of bio-agents is feasible for threat-level concentrations in practical battlefield environments at sufficient ranges to provide for early warning.